# Quantitative Magnetohydrodynamic Modelling of Flux Pumping in ASDEX Upgrade

### v1

@@ Abstract:
The sawtooth-free hybrid scenario has been achieved recently in ASDEX Upgrade (AUG) with applied non-inductive current sources and auxiliary heating [A. Burckhart et al 2023 Nucl. Fusion 63 126056]. Control experiments in AUG suggest that the self-regulating magnetic flux pumping mechanism, characterized by anomalous current redistribution, is responsible for clamping the central safety factor () close to unity, thereby preventing the sawtooth onset. This work presents a numerical and theoretical investigation of flux pumping in the AUG hybrid scenario based on the two-temperature, visco-resistive, full magnetohydrodynamic (MHD) model with the JOREK code. To quantitatively model the flux pumping, we choose realistic parameters, plasma configurations, and source terms based on AUG experiments. During the initial saturation stage of the unstable 1/1 quasi-interchange mode (on millisecond timescales), exhibits fast under-damped oscillation and reaches a value closer to unity, which is attributed to the self-regulation of core plasma and the fast dynamo effect on the order of V/m. On the longer resistive diffusion timescale of seconds, the slow negative dynamo effect on the order of mV/m induced by the 1/1 MHD instability plays an effective role in flux pumping, which provides quantitative agreement with experimental observations for the first time. The final saturated 1/1 MHD instability exhibits features of the quasi-interchange mode and tearing mode, and the associated convective plasma flow velocity is a few m/s. The toroidal negative electric field from the slow dynamo dominantly offsets the positive current drive and continuously redistributes the current density and pressure. As a result, is maintained close to unity due to the low-shear profiles of current density and pressure in the plasma core, and the system enters into a sawtooth-free and quasi-stationary helical state.
@@ 1 Research background
As a first quantitative milestone, this work investigates the flux pumping experiments in AUG using the JOREK code[1, 34]. To accurately model the nonlinear plasma dynamics including the MHD dynamo effect in the AUG flux pumping discharge [17], we adopt reconstructed experimental profiles and realistic parameters. Based on the two-temperature, visco-resistive, full MHD model, simulations have been run over a resistive diffusion timescale of seconds. Two self-regulating stages of core plasma are observed, respectively caused by the fast and slow dynamo effects [24]. Firstly, at the initial saturation stage of the unstable 1/1 quasi-interchange mode (on millisecond timescales), shows rapid and under-damped oscillation at 5.8 kHz, and the system self-regulates into a state where is closer to unity. The oscillation of is mainly caused by the fast dynamo effect on the order of V/m, which acts much faster than resistive diffusion and is mainly generated by the quasi-interchange. Secondly, further simulations on the resistive diffusion timescale of seconds verify the effective role of the slow negative dynamo effect in flux pumping, where the associated 1/1 mode exhibits the features of both ideal quasi-interchange and resistive tearing and the convective plasma flow velocity is a few m/s. The amplitude of the obtained negative toroidal dynamo electric field ( mV/m) is comparable in magnitude to the experimental observations, while previous qualitative results show substantial overestimation, primarily attributed to the choice of increased resistivity [30, 17]. Especially, the slow negative dynamo can self-regulate its profile and amplitude within a certain range to counteract the positive current drive and ensure the continuous redistribution of current density and pressure in the plasma core, thereby keeping close to unity and leading to a sawtooth-free quasi-stationary helical state.
@@ 2.1 A brief review of the AUG experiment
JOREK [35, 36] simulations are carried out to investigate the flux pumping phases observed in the AUG discharge #36663 [17]. This discharge is mainly designed to experimentally study the transition behaviour between sawtoothing and flux pumping phases by adjusting the intensities of NBI heating and central co-ECCD, as shown by the schematic in Fig. 1. The detailed time traces of this discharge can be found in Fig. 3 of Ref. [17].
@@
The repeated transitions between sawtoothing and flux pumping phases clearly confirm previous theoretical prediction on the threshold of to suppress sawteeth through flux pumping, which is found to result from the dynamo effect by the pressure-gradient driven 1/1 quasi-interchange mode[30]. However, the nature of the 1/1 mode and its role in the current redistribution observed during the flux pumping phases of the AUG discharge remain unclear [17]. The JOREK simulations of flux pumping for phase III of AUG discharge #36663 presented in this work aim to provide an intuitive understanding of the underlying dynamo effect and plasma self-regulation.
@@ 2.2 Two-temperature full MHD model
In previous flux pumping simulations, two dominant mechanisms responsible for the current redistribution are identified [30]. The first is the negative dynamo associated with the convection flow and magnetic field distortion of a 1/1 quasi-interchange mode, which counteracts the peaking tendency of current density in a driven system. The second mechanism is the flattening and lifting of resistivity caused by the flattened electron temperature profile, which strengthens the resistive current diffusion in the plasma core. Since the initial electron temperature profile and its subsequent evolution in the presence of heating sources could deviate from ions, separated temperature equations are adopted for electrons and ions in JOREK simulations. Meanwhile, in simulations of the quasi-interchange mode driven by pressure gradient, parallel magnetic field perturbation is usually comparable with the normalized pressure perturbation (). Therefore, the full MHD model is required to describe the self-consistent evolution of and obtain the correct linear properties of the instability [34]. The diamagnetic terms are not considered in the present simulations due to the significant challenge of long timescale simulations [34], which will be overcome in future work on the development of the extended MHD model. The following normalized, two-temperature, single-fluid, visco-resistive, full MHD model is adopted in the base cylindrical coordinate system
where the magnetic field , the plasma current density , and the total pressure . is the magnetic vector potential, is the electrostatic potential, is the plasma velocity, is the mass density, is the ion (electron) pressure, is the ion (electron) temperature, is an axisymmetric equilibrium function of the poloidal magnetic flux , defined to satisfy the Grad–Shafranov equilibrium (the equilibrium toroidal magnetic field ) and is kept constant over time, is the current source in the toroidal direction (), is the particle source, is the ion (electron) heating source, is the ratio of specific heats of a monatomic gas, is the resistivity (applied only in the toroidal direction), is the dynamic viscosity, is the perpendicular (parallel) particle diffusivity, is the perpendicular (parallel) heat conductivity.
@@ 2.4 Source terms and diffusion coefficients
We choose the realistic Spitzer resistivity and a semi-empirical viscosity (with the value between the Braginskii perpendicular and gyro viscosities) [39], both depending on the electron temperature, i.e., and . and are the values of resistivity and viscosity on the axis. Unless otherwise specified, we use for both 2D and 3D simulations. And (kinematic viscosity ). The corresponding magnetic Prandtl number is , the Hartmann number is 222Note that due to the low resistivity and viscosity, the Hartmann number in this work is much larger than that in previous MHD simulations for tokamak and RFP plasmas [40, 32, 30, 41, 25, 26]. As will be discussed in Subsec. 6.2, the preliminary parameter scan shows that increasing the Hartmann number results in the transition of core plasma from flux pumping towards sawtooth-like oscillation and saturated resistive internal kink mode. Detailed simulation results of the viscosity and resistivity dependence of different plasma states will be reported in a separate paper., and the Lundquist number is (characteristic length ) [41]. Realistic temperature-dependent parallel heat conductivities are respectively implemented for ions and electrons, i.e., . is calculated based on the ion (electron) temperature on the axis using Spitzer-Haerm formula [42] (at high temperature, the parallel heat conductivities are typically overestimated, and they are reduced by a factor of 30 due to the heat flux limit correction[43, 44]), where and . The radial profiles of perpendicular heat conductivities are calculated from the axisymmetric stationary conditions for ion and electron temperatures, i.e., satisfying . As a result, the temperature amplitudes and profiles can be roughly maintained in the 2D simulations without MHD instabilities. The values on the axis are and , respectively. The anisotropy of electron heat conductivity () is of the order of . For the continuity equation as Eq. 3, the perpendicular particle diffusivity is set as 0, while a constant parallel particle diffusivity is adopted to average density perturbations along the magnetic field line and avoid negative density values. The detailed values and distributions of the adopted parameters are also listed in Table 1.
@@
For the numerical discretization in JOREK, the implicit Gear scheme for time advancement, the G1-continuous finite element (quadrangular bi-cubic Bezier elements) based on the weak form for the () poloidal plane and a finite Fourier series expansion in the toroidal direction () are adopted [38]. The time step of and a polar grid with the resolution of are used. Parameter scans in both time step and spatial resolution were conducted to ensure converged linear growth rates and saturation amplitudes of the 1/1 quasi-interchange. A convergence scan of different truncations on the Fourier series of toroidal modes shows that the 3D simulation with only n = 0 and 1 harmonics cannot correctly capture the nonlinear evolutions of the equilibrium and the n = 1 mode. Therefore, the maximum toroidal mode number is set to 4 in the 3D simulations to guarantee convergence.
@@
| Spitzer resistivity | ||
| Dynamical viscosity | ||
@@
| Parallel heat conductivity (e) | ||
| Parallel heat conductivity (i) | ||
| Perpendicular heat conductivity (e) | ||
| Perpendicular heat conductivity (i) | ||
| Parallel particle diffusivity | uniform | |
@@
| Magnetic Prandtl number [] | ||
| Hartmann number [] | () | |
| Lundquist number () | () | |
@@
| T | ||
| 1.04 | ||
@@ 3.2 Linear characteristics of the MHD instability in the 3D simulation
In this subsection, we first analyze the linear characteristics of the MHD instability for the equilibrium shown in Fig. 2. In the 3D simulation, toroidal harmonics are limited to n 4, and the toroidal mode number of the most unstable instability is n = 1. The time trace of normalized magnetic and kinetic energies is plotted in Fig. 6 (a). During 0.25 1.3 ms, the n = 1 mode grows linearly with a growth rate of 19.46 ms-1, while the other modes with n = 2 4 are not growing linearly, but are excited by the n = 1 via nonlinear mode coupling. The dominant poloidal mode number is m = 1. In Fig. 6 (b), the linear eigenmode structure of the n = 1 instability is plotted in the form of the poloidal flow field. Within the central low magnetic shear region, a strong convection cell forms. The plasma flow exhibits a large amplitude in the whole plasma core, which aligns with the feature of the 1/1 quasi-interchange mode and is quite distinct from that of the 1/1 internal kink mode. In the latter case, the poloidal flow component dominates around the rational surface. A more detailed comparison of the linear eigenmode structures of 1/1 quasi-interchange mode and 1/1 internal kink mode is presented in Fig. 14 of Ref. [30].
@@ 3.3 Self-regulation and fast dynamo in the early saturation stage
Before we investigate the flux pumping occurring in the resistive diffusion timescale, a rapid oscillation of due to the fast dynamo effect in the 3D simulation is noteworthy when the 1/1 quasi-interchange mode becomes prominent in its early saturation phase, as shown in Fig. 7 (a). At about 1.3 ms, first drops rapidly by 0.03 on a short timescale of 0.1 ms. Then oscillates around 1.013 with the frequency of 5.8 kHz for a few milliseconds until it becomes relatively stable. After the fast oscillation phase in Fig. 7 (a), enters a slow decay phase with a characteristic timescale of hundreds of milliseconds, as plotted by the solid line in Fig. 9 (c). It will be shown in Sec. 4 that the ultimate saturation of is mainly determined by the balance between the current peaking (due to current drive and auxiliary heating) and the negative slow dynamo effect (contributed by non-axisymmetric fluctuations of plasma velocity and magnetic field ), which occurs on the resistive diffusion timescale, much slower than the oscillation of observed here.
@@
The rapidly damped oscillation of and dynamo emf is essentially a self-regulating relaxation of the core plasma, which can potentially enhance the robustness of flux pumping discharge. In tokamak experiments, after the initial current ramp-up, the discharge will likely begin with a slight deviation of from the target value for maintaining the quasi-stationary flux pumping state. However, the slight error in equilibrium will be roughly reduced by the fast dynamo effect on the short saturation timescale of 1/1 quasi-interchange. Subsequently, the equilibrium will be further refined by the competition between current-driven mechanisms and the slow dynamo effect on the resistive diffusion timescale. In such a case, to avoid the sawtooth onset via flux pumping in experiments, it is not necessary to manually fine-tune the central current and pressure profiles, but the proper equilibrium with close to unity can be achieved by the self-regulation of the core plasma.

### v3

@@ Abstract:
The sawtooth-free hybrid scenario has been achieved recently in ASDEX Upgrade (AUG) with applied non-inductive current sources and auxiliary heating [A. Burckhart et al 2023 Nucl. Fusion 63 126056]. Control experiments in AUG suggest that the self-regulating magnetic flux pumping mechanism, characterized by anomalous current redistribution, is responsible for clamping the central safety factor () close to unity, thereby preventing the sawtooth onset. This work presents a numerical and theoretical investigation of flux pumping in the AUG hybrid scenario based on the two-temperature, visco-resistive, full magnetohydrodynamic (MHD) model with the JOREK code. To quantitatively model the flux pumping, we choose realistic parameters, plasma configurations, and source terms based on AUG experiments. During the initial saturation stage of the unstable 1/1 quasi-interchange mode (on millisecond timescales, i.e., thousands of Alfvén times), exhibits fast under-damped oscillation and reaches a value closer to unity, which is attributed to the self-regulation of core plasma and the fast dynamo effect on the order of V/m. On the longer resistive diffusion timescale of seconds (millions of Alfvén times), the slow negative dynamo effect on the order of mV/m induced by the 1/1 MHD instability plays an effective role in flux pumping, which provides quantitative agreement with experimental observations for the first time. The final saturated 1/1 MHD instability exhibits features of the quasi-interchange mode and tearing mode, and the associated convective plasma flow velocity is a few m/s. The toroidal negative electric field from the slow dynamo dominantly offsets the positive current drive and continuously redistributes the current density and pressure. As a result, is maintained close to unity due to the low-shear profiles of current density and pressure in the plasma core, and the system enters into a sawtooth-free and quasi-stationary helical state.
@@ 1 Research background
As a first quantitative milestone, this work investigates the flux pumping experiments in AUG using the JOREK code[1, 34]. To accurately model the nonlinear plasma dynamics including the MHD dynamo effect in the AUG flux pumping discharge [17], we adopt reconstructed experimental profiles and realistic parameters. Based on the two-temperature, visco-resistive, full MHD model, simulations have been run over a resistive diffusion timescale of seconds. Two self-regulating stages of core plasma are observed, respectively caused by the fast and slow dynamo effects [24]. Firstly, at the initial saturation stage of the unstable 1/1 quasi-interchange mode (on millisecond timescales, i.e., thousands of Alfvén times), shows rapid and under-damped oscillation at 5.8 kHz, and the system self-regulates into a state where is closer to unity. The oscillation of is mainly caused by the fast dynamo effect on the order of V/m, which acts much faster than resistive diffusion and is mainly generated by the quasi-interchange. Secondly, further simulations on the resistive diffusion timescale of seconds (millions of Alfvén times) verify the effective role of the slow negative dynamo effect in flux pumping, where the associated 1/1 mode exhibits the features of both ideal quasi-interchange and resistive tearing, and the convective plasma flow velocity is a few m/s. The amplitude of the obtained negative toroidal dynamo electric field ( mV/m) is comparable in magnitude to the experimental observations, while previous qualitative results show substantial overestimation, primarily attributed to the choice of increased resistivity [30, 17]. Especially, the slow negative dynamo can self-regulate its profile and amplitude within a certain range to counteract the positive current drive and ensure the continuous redistribution of current density and pressure in the plasma core, thereby keeping close to unity and leading to a sawtooth-free quasi-stationary helical state.
@@ 2.1 A brief review of the AUG experiment
JOREK [35, 36] simulations are carried out to investigate the flux pumping phases observed in the AUG discharge #36663 [17]. This discharge is mainly designed to experimentally study the transition behaviour between sawtoothing and flux pumping phases by adjusting the intensities of NBI heating and central co-ECCD, as shown by the schematic in Fig. 1. The detailed time traces of this discharge can be found in Fig. 3 of Ref. [17]. Note that the horizontal and vertical axes of the schematic only represent the relative changes in the strength of co-ECCD and NBI, respectively, and do not represent any actual values. The schematic also does not indicate any plasma behaviour in the absence of co-ECCD and NBI.
@@
The repeated transitions between sawtoothing and flux pumping phases clearly confirm previous theoretical prediction on the threshold of to suppress sawteeth through flux pumping, which is found to result from the dynamo effect by the pressure gradient driven 1/1 quasi-interchange mode[30]. However, the nature of the 1/1 mode and its role in the current redistribution observed during the flux pumping phases of the AUG discharge remain unclear [17]. The JOREK simulations of flux pumping for phase III of AUG discharge #36663 presented in this work aim to provide an intuitive understanding of the underlying dynamo effect and plasma self-regulation.
@@ 2.2 Two-temperature full MHD model
In previous flux pumping simulations, two dominant mechanisms responsible for the current redistribution are identified [30]. The first is the negative dynamo associated with the convection flow and magnetic field distortion of a 1/1 quasi-interchange mode, which counteracts the peaking tendency of current density in a driven system. The second mechanism is the flattening and lifting of resistivity caused by the flattened electron temperature profile, which strengthens the resistive current diffusion in the plasma core. Since the initial electron temperature profile and its subsequent evolution in the presence of heating sources could deviate from ions, separated temperature equations are adopted for electrons and ions in JOREK simulations. Nevertheless, the Ohmic heating term (much less than the auxiliary heating power) and the equipartition terms between the temperatures of two species (causing numerical instabilities) have not been considered in this work. Meanwhile, in simulations of the quasi-interchange mode driven by pressure gradient, parallel magnetic field perturbation is usually comparable with the normalized pressure perturbation (). Therefore, the full MHD model is required to describe the self-consistent evolution of and obtain the correct linear properties of the instability [34]. The Hall term and the diamagnetic terms are not considered in the present simulations due to the significant challenge of long timescale simulations [34], which will be overcome in future work on the development of the extended MHD model. The following normalized, two-temperature, single-fluid, visco-resistive, full MHD model is adopted in the base cylindrical coordinate system
where the magnetic field , the plasma current density , and the total pressure . is the magnetic vector potential, is the electrostatic potential, is the plasma velocity, is the mass density, is the ion (electron) pressure, is the ion (electron) temperature, is an axisymmetric equilibrium function of the poloidal magnetic flux , defined to satisfy the Grad–Shafranov equilibrium (the equilibrium toroidal magnetic field ) and is kept constant over time, is the current source in the toroidal direction (), is the particle source, is the ion (electron) heating source, is the ratio of specific heats of a monatomic gas, is the resistivity (applied only in the toroidal direction to improve the numerical stability), is the dynamic viscosity, is the perpendicular (parallel) particle diffusivity, is the perpendicular (parallel) heat conductivity.
@@ 2.4 Source terms and diffusion coefficients
We choose the realistic Spitzer resistivity and a semi-empirical viscosity (with the value between the Braginskii perpendicular and gyro viscosities) [39], both depending on the electron temperature, i.e., and . and are the values of resistivity and viscosity on the axis. Unless otherwise specified, we use for both 2D and 3D simulations. And (kinematic viscosity ). The corresponding magnetic Prandtl number is , the Hartmann number is 222Note that due to the low resistivity and viscosity, the Hartmann number in this work is much larger than that in previous MHD simulations for tokamak and RFP plasmas [40, 32, 30, 41, 25, 26]. As will be discussed in Subsec. 6.2, the preliminary parameter scan shows that increasing the Hartmann number results in the transition of core plasma from flux pumping towards sawtooth-like oscillation and saturated resistive internal kink mode. Detailed simulation results of the viscosity and resistivity dependence of different plasma states will be reported in a separate paper., and the Lundquist number is (characteristic length ) [41]. Realistic temperature-dependent parallel heat conductivities are respectively implemented for ions and electrons, i.e., . is calculated based on the ion (electron) temperature on the axis using Spitzer-Haerm formula [42] (at high temperature, the parallel heat conductivities are typically overestimated, and they are reduced by a factor of 30 due to the heat flux limit correction[43, 44]), where and . The radial profiles of perpendicular heat conductivities are calculated from the axisymmetric stationary conditions for ion and electron temperatures, i.e., satisfying . As a result, the temperature amplitudes and profiles can be roughly maintained in the 2D simulations without MHD instabilities. The values on the axis are and , respectively. The anisotropy of electron heat conductivity () is of the order of . In the continuity equation as Eq. 3, the perpendicular particle diffusivity is set as 0, and in order to maintain numerical stability, a constant parallel particle diffusivity is adopted to average density perturbations along the magnetic field line and avoid negative density values. The detailed values and distributions of the adopted parameters are also listed in Table 1.
@@
In JOREK, the implicit Gear scheme for time advancement, the G1-continuous finite element (quadrangular bi-cubic Bezier elements) [45, 36, 46] based on the weak form for the () poloidal plane and a finite Fourier series expansion in the toroidal direction () [38] are adopted. The time step of and a polar grid with the resolution of are used. Parameter scans in both time step and spatial resolution were conducted to ensure converged linear growth rates and saturation amplitudes of the 1/1 quasi-interchange. A convergence scan of different truncations on the Fourier series of toroidal modes shows that the 3D simulation with only n 1 harmonics cannot correctly capture the flux pumping. With n 2 and n 4, the flux pumping can be obtained in both cases, with the relative differences in the saturated and toroidal current density of the order of 1%. Therefore, the maximum toroidal mode number is set to 4 in the 3D simulations to guarantee convergence of the saturated solution under acceptable computational resources and numerical uncertainties.
@@
Spitzer resistivity |
(JOREK units) |
Dynamical viscosity |
(JOREK units) |
@@
Parallel heat conductivity (e) |
(JOREK units) |
Parallel heat conductivity (i) |
(JOREK units) |
Perpendicular heat conductivity (e) |
(JOREK units) |
Perpendicular heat conductivity (i) |
(JOREK units) |
Parallel particle diffusivity |
(JOREK units) |
uniform (for the sake of numerical stability) |
@@
Magnetic Prandtl number
Hartmann number
Lundquist number
@@
Central magnetic field () | T | |
Central safety factor () | 1.04 | |
@@ 3.2 Linear characteristics of the MHD instability in the 3D simulation
In this subsection, we first analyze the linear characteristics of the MHD instability for the equilibrium shown in Fig. 2. In the 3D simulation, toroidal harmonics are limited to n 4, and the toroidal mode number of the most unstable instability is n = 1. The time trace of normalized magnetic and kinetic energies is plotted in Fig. 6 (a). During 0.25 1.3 ms, the n = 1 mode grows linearly with a growth rate of 19.46 ms-1, while the other modes with n = 2 4 are not growing linearly, but are excited by the n = 1 via nonlinear mode coupling. The dominant poloidal mode number is m = 1. In Fig. 6 (b), the linear eigenmode structure of the n = 1 instability is plotted in the form of the poloidal flow field. Within the central low magnetic shear region, a strong convection cell forms. The plasma flow exhibits a large amplitude in the whole plasma core, which aligns with the feature of the 1/1 quasi-interchange mode and is quite distinct from that of the 1/1 internal kink mode. In the latter case, the poloidal flow dominates around the rational surface. A more detailed comparison of the linear eigenmode structures of 1/1 quasi-interchange mode and 1/1 internal kink mode is presented in Fig. 14 of Ref. [30].
@@ 3.3 Self-regulation and fast dynamo in the early saturation stage
Before we investigate the flux pumping occurring in the resistive diffusion timescale, the rapid oscillations of 333Note that the profiles shown in this paper are mainly obtained by tracing the axisymmetric (n = 0) component of the magnetic field, with the exception of Fig. 12, which is calculated by tracing the total magnetic field and considering the helical magnetic axis as reference [47]. and the central plasma pressure () due to the fast dynamo effect in the 3D simulation are noteworthy when the 1/1 quasi-interchange mode becomes prominent in its early saturation phase, as shown in Fig. 7 (a). Taking the evolution of as an example, at about 1.3 ms, first drops rapidly by 0.03 on a short timescale of 0.1 ms. This is much faster than the evolution of in the 2D case that occurs in the resistive diffusion timescale. Then oscillates around 1.013 with the frequency of 5.8 kHz for a few milliseconds until it becomes relatively stable. After the fast oscillation phase in Fig. 7 (a), enters a slow decay phase with a characteristic timescale of hundreds of milliseconds, as plotted by the solid line in Fig. 9 (c). It will be shown in Sec. 4 that the ultimate saturation of is mainly determined by the balance between the current peaking (due to current drive and auxiliary heating) and the negative slow dynamo effect (contributed by non-axisymmetric fluctuations of plasma velocity and magnetic field ), which occurs on the resistive diffusion timescale, much slower than the oscillation of observed here.
@@
The rapidly damped oscillation of , , and dynamo emf is essentially a self-regulating relaxation of the core plasma, which can potentially enhance the robustness of flux pumping discharge. In tokamak experiments, after the initial current ramp-up, the discharge will likely begin with a slight deviation of from the target value for maintaining the quasi-stationary flux pumping state. However, the slight error in equilibrium will be roughly reduced by the fast dynamo effect on the short saturation timescale of 1/1 quasi-interchange. Subsequently, the equilibrium will be further refined by the competition between current-driven mechanisms and the slow dynamo effect on the resistive diffusion timescale. In such a case, to avoid the sawtooth onset via flux pumping in experiments, it is not necessary to manually fine-tune the central current and pressure profiles, but the proper equilibrium with close to unity can be achieved by the self-regulation of the core plasma.